# Thermodynamic Complex Phase Diagram for the Born-Infeld AdS Black Holes with Reentrant Phase Transition

### v1

@@ Abstract
The Lee-Yang phase transition theory applied in the anti-de Sitter (AdS) black hole has inspired the exploration of complex phase diagram and supercritical phenomena in black hole thermodynamics. In this study, we extend the approach to the four dimensional Born-Infeld AdS black hole. This system exhibits a rich phase structure, including reentrant phase transitions, due to the modulation of the Born-Infeld nonlinear parameter. Through the analysis by Lee-Yang zeros, we obtained the complex phase diagram of the Born-Infeld AdS black hole and derived the supercritical crossover line—Widom line. It strictly originates from the first-order stable critical point. The results indicate that Born-Infeld nonlinear effects significantly alter the types and characteristics of phase transition in critical region, while do not disrupt the uniqueness of the Widom line in supercritical region. Our study uncovers a universal simplified feature of the thermodynamic behavior of nonlinear gravitational systems in supercritical region. It also deepens our understanding of the fundamental connection between critical phenomena and continuous phase transitions in the extended phase space of black holes.
@@ I Introduction
With the deepening of research into black hole thermodynamics, people have discovered a fascinating duality between gravitational systems and classical fluids Maldacena:1997re ; Chamblin1999 ; Kastor2009 ; Dolan2011 ; Kubiznak:2012wp ; Kubiznak:2016qmn ; Mann2025 . In the extended phase space, a negative cosmological constant is assigned a physical meaning as the thermodynamic pressure, with its conjugate quantity corresponding to the thermodynamic volume. Based on this framework, the charged AdS black hole has been shown to exhibit phase transition behavior highly similar to that of van der Waals fluids Kubiznak:2012wp ; Wei2022a ; Wei2023 . Even more excitingly, black hole systems also display reentrant phase transitions: as the temperature decreases under constant pressure, the black hole system can undergo a sequential phase transition from “large black hole (LBH) $\to$ small black hole (SBH) $\to$ large black hole (LBH)”. Although this phenomenon was initially discovered in higher-dimensional ($d\geq 6$) rotating AdS black holes and in Lovelock gravity theories, the introduction of Born-Infeld (BI) nonlinear electrodynamics provides a new mechanism to simulate such rich behavior in four-dimensional spacetime Altamirano:2013ane ; Frassino:2014pha ; Gunasekaran:2012dq . The BI nonlinear term breaks the linear constraints of Maxwellian electrodynamics Altamirano:2014tva ; Breton:2004qa and this nonlinear effect plays a thermodynamic role similar to high-dimensional gravitational corrections, enabling four-dimensional BI-AdS black holes to exhibit phase structures highly sensitive to the parameter $b$ —including no phase transition, reentrant phase transitions, and standard first-order phase transitions Chemissany:2008fy ; Banerjee:2011cz ; Banerjee:2012zm ; Miskovic:2008ck ; Fernando:2006gh ; Dey:2004yt ; Zou:2013owa . This makes BI-AdS black hole an ideal theoretical model for exploring the impact of nonlinear effects on the thermodynamic structure of spacetime.
@@ II.1 Thermodynamics
In the AdS spacetime, the action of the Einstein-Born-Infeld theory can be written as Miskovic:2008ck ; Fernando:2006gh
@@
where the thermodynamic volume $V=4\pi r_{h}^{3}/3$ and other related thermodynamic quantities are
$M$ $=\frac{r_{h}}{2}+\frac{4}{3}\pi Pr_{h}^{3}+\frac{b^{2}r_{h}^{3}}{3}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)+\frac{2Q^{2}}{3r_{h}^{2}}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right),$ (8)
$T$ $=\frac{1}{4\pi r_{h}}\left(1+8\pi Pr_{h}^{2}+2b^{2}r_{h}^{2}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)\right),$ (9)
@@
$B$ $=\frac{2}{3}br_{h}^{3}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}+\frac{Q^{2}}{3br_{h}}\right)\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right).$ (12)
According to the standard definition of thermodynamics, Gibbs free energy and heat capacity at constant pressure can be obtained
$G$ $=\frac{1}{4}r_{h}-\frac{2\pi}{3}Pr_{h}^{3}-\frac{b^{2}r_{h}^{3}}{6}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)+\frac{2Q^{2}}{3r_{h}}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right),$ (13)
$C_{P}$ $=\frac{2\pi r_{h}^{4}\left[1-2\left(4Pr_{h}^{2}+b^{2}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)\right)\right]}{2Q^{2}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)r_{h}^{2}+b^{2}\left(1-\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}\right)^{2}r_{h}^{2}+4Pr_{h}^{2}\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}}.$ (14)
@@ II.2 Phase Structure
which results that
$T_{\rm c}=\frac{1-8x}{\pi v_{\rm c}},\quad v_{\rm c}=\left(\frac{1}{x_{k}^{2}}-\frac{16}{b^{2}}\right)^{\frac{1}{4}},\quad P_{\rm c}=\frac{1-16x}{2\pi v_{\rm c}^{2}}-\frac{b^{2}}{4\pi}\left(1-\frac{1}{v_{\rm c}^{2}x}\right)$ (18)
where the parameters $x_{k}$, $p$ and $q$ are given by
$x_{k}$ $=2\sqrt{-\frac{p}{3}}\cos\left(\frac{1}{3}\arccos\left(\frac{3q}{2p}\sqrt{-\frac{3}{p}}\right)-\frac{2\pi k}{3}\right)\quad(k=0,1,2),$ (19)
$p$ $=-\frac{3b^{2}}{32},\quad q=\frac{b^{2}}{256}.$ (20)
For $x_{2}$ (i.e., $k=2$), the values of the corresponding critical point are complex, so the system can have at most two critical points. According to reference Altamirano:2014tva ; Xu:2019yub , the parameter $b$ can be divided into three cases, corresponding to the no-phase-transition region, the reentrant phase-transition region, and the standard first-order phase-transition region.
Non-phase transition regime ($b<0.3536$)  In this parameter range, the system exhibits the simplest thermodynamic behavior. It is analogous to that of the Schwarzschild-AdS black hole, with no critical points or phase transition phenomena present.
Reentrant phase transition regime ($0.3536<b<0.5$)  In this parameter range, the system has two critical points and exhibits reentrant phase transition phenomena. As shown in the left diagram in Fig. 1, when the pressure is fixed at $P=0.98P_{z}$ ($P_{z}$ corresponds to the pressure at the onset of the zeroth-order phase transition), as the temperature decreases to $T_{1}$, the system undergoes a first-order phase transition from large black hole to small black hole. The Gibbs free energy of the two phases is continuous, but its first derivative is discontinuous, corresponding to the characteristics of a standard first-order phase transition. This phase transition process is depicted by a orange solid line in the right diagram in Fig. 1, showing a typical gas-liquid-type coexistence curve and ending at a critical point 1 $(P_{\rm c}^{(1)},T_{\rm c}^{(1)})$, where the heat capacity $C_{P}$ and isothermal compressibility $\kappa_{T}$ exhibit divergent behavior. When the temperature is further lowered to $T_{0}<T_{1}$, the system undergoes a transition from small black hole to large black hole, with the Gibbs free energy exhibiting a discontinuous jump $\Delta G\neq 0$ at temperature $T_{0}$, marking the occurrence of a zeroth-order phase transition. This zeroth-order phase transition starts at the point $(P_{z},T_{z})$ and ends at the triple point $(P_{t},T_{t})$. This phase transition region is extremely narrow, and we have determined its corresponding red solid line in the right diagram in Fig. 1. It is worth noting that this red line is not a traditional phase coexistence curve. Along this line, small black holes have a lower Gibbs free energy and remain in a locally stable state rather than coexisting phases; accordingly, the endpoint of this curve does not satisfy the critical point criterion. In a word, the first-order phase transition at high temperatures and the zeroth-order phase transition at low temperatures together form a complete reentrant phase diagram, profoundly reflecting the nontrivial coupling effects of BI nonlinear electrodynamics with the AdS background Rasheed:1997ns .
@@
Standard first-order phase transition regime ($b>0.5$)  In this parameter range, nonlinear effects become very weak, and the system behavior approaches that of linear Maxwell theory. As shown in Fig. 2, the system has only one physical critical point. The the left diagram in Fig. 2 exhibits standard swallowtail behavior when the temperature decreases to $T_{2}$, the black hole undergoes a phase transition from small black hole to large black hole, displaying the typical first-order phase transition behavior analogous to that of four-dimensional RN-AdS black holes Kubiznak:2012wp .
@@ III Complex Phase diagram and Widom line
$2-2b^{2}r_{h}^{4}+\left(-r_{h}^{2}+2b^{2}r_{h}^{4}+8\pi Pr_{h}^{4}\right)\sqrt{1+\frac{1}{b^{2}r_{h}^{4}}}=0.$ (21)

### v2

# Complex Plane Phase Diagram and Widom Line for the Born-Infeld Black Holes with Reentrant Phase Transition

@@ Abstract
The Lee-Yang phase transition theory applied in the anti-de Sitter (AdS) black hole has inspired the exploration of complex phase diagram and supercritical phenomena in black hole thermodynamics. In this study, we extend the approach to the four dimensional Born-Infeld AdS black hole. This system exhibits a rich phase structure, including reentrant phase transitions, due to the modulation of the Born-Infeld nonlinear parameter. Through the Lee-Yang zeros, we obtained the complex phase diagram of the Born-Infeld AdS black hole and derived the supercritical crossover line—Widom line, which strictly originates from the first-order stable critical point. The results indicate that Born-Infeld nonlinear effects significantly alter the types and characteristics of phase transition in critical region, while do not disrupt the uniqueness of the Widom line in supercritical region. Our study uncovers a universal simplified feature of the thermodynamic behavior of nonlinear gravitational systems in supercritical region. It also deepens our understanding of the fundamental connection between critical phenomena and continuous phase transitions in the extended phase space of black holes.
@@ I Introduction
With the deepening of research into black hole thermodynamics, people have discovered a fascinating duality between gravitational systems and classical fluids Maldacena:1997re ; Chamblin1999 ; Kastor2009 ; Dolan2011 ; Kubiznak:2012wp ; Kubiznak:2016qmn ; Mann2025 . In the extended phase space, a negative cosmological constant is assigned a physical meaning as the thermodynamic pressure, with its conjugate quantity corresponding to the thermodynamic volume. Based on this framework, the charged AdS black hole has been shown to exhibit phase transition behavior highly similar to that of van der Waals fluids Kubiznak:2012wp ; Kubiznak:2016qmn ; Mann2025 ; Wei2022a ; Wei2023 ; Karch2015 ; Mancilla2025 . Even more excitingly, black hole systems also display reentrant phase transitions: as the temperature decreases under constant pressure, the black hole system can undergo a sequential phase transition from “large black hole (LBH) $\to$ small black hole (SBH) $\to$ large black hole (LBH)”. Although this phenomenon was initially discovered in higher-dimensional ($d\geq 6$) rotating AdS black holes and in Lovelock gravity theories, the introduction of Born-Infeld (BI) nonlinear electrodynamics provides a new mechanism to simulate such rich behavior in four-dimensional spacetime Altamirano:2013ane ; Frassino:2014pha ; Gunasekaran:2012dq . The BI nonlinear term breaks the linear constraints of Maxwellian electrodynamics Altamirano:2014tva ; Breton:2004qa and this nonlinear effect plays a thermodynamic role similar to high-dimensional gravitational corrections, enabling four-dimensional BI-AdS black holes to exhibit phase structures highly sensitive to the parameter $b$ —including no phase transition, reentrant phase transitions, and standard first-order phase transitions Chemissany:2008fy ; Banerjee:2011cz ; Banerjee:2012zm ; Miskovic:2008ck ; Fernando:2006gh ; HENDI2015157 ; Hendi2016a ; Hendi2016b ; Aghbolagh2025 ; Dey:2004yt ; Zou:2013owa . This makes BI-AdS black hole an ideal theoretical model for exploring the impact of nonlinear effects on the thermodynamic structure of spacetime.
@@ II.1 Thermodynamics
In the AdS spacetime, the action of the Einstein-Born-Infeld theory can be written as Miskovic:2008ck ; Fernando:2006gh ; HENDI2015157 ; Hendi2016a ; Hendi2016b ; Aghbolagh2025
@@
where the thermodynamic volume $V=4\pi r_{h}^{3}/3$ and other related thermodynamic quantities are (to simplify the expression, we replace $\sqrt{1+\frac{Q^{2}}{b^{2}r_{h}^{4}}}$ with the letter $A$)
$M$ $=\frac{r_{h}}{2}+\frac{4}{3}\pi Pr_{h}^{3}+\frac{b^{2}r_{h}^{3}}{3}\left(1-A\right)+\frac{2Q^{2}}{3r_{h}^{2}}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right),$ (8)
$T$ $=\frac{1}{4\pi r_{h}}\left(1+8\pi Pr_{h}^{2}+2b^{2}r_{h}^{2}\left(1-A\right)\right),$ (9)
@@
$B$ $=\frac{2}{3}br_{h}^{3}\left(1-A\right)+\frac{Q^{2}}{3br_{h}}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right).$ (12)
According to the standard definition of thermodynamics, the Gibbs free energy and heat capacity at constant pressure can be obtained
$G$ $=\frac{1}{4}r_{h}-\frac{2\pi}{3}Pr_{h}^{3}-\frac{b^{2}r_{h}^{3}}{6}\left(1-A\right)+\frac{2Q^{2}}{3r_{h}}\,_{2}F_{1}\left(\frac{1}{4},\frac{1}{2};\frac{5}{4};-\frac{Q^{2}}{b^{2}r_{h}^{4}}\right),$ (13)
$C_{P}$ $=\frac{2\pi Ar_{h}^{4}-4\pi Q^{2}r_{h}^{2}+16\pi^{2}PAr_{h}^{6}+4\pi b^{2}Ar_{h}^{6}-4\pi b^{2}r_{h}^{6}}{2Q^{2}-2b^{2}r_{h}^{4}-Ar_{h}^{2}+2b^{2}Ar_{h}^{4}+8P\pi Ar_{h}^{4}}.$ (14)
@@ II.2 Phase Structure
According to reference Altamirano:2014tva ; Xu:2019yub , the parameter $b$ can be divided into three cases, corresponding to the no-phase-transition region, the reentrant phase-transition region, and the standard first-order phase-transition region.
Non-phase transition regime ($b<0.3536$)  In this parameter range,the system exhibits the simplest thermodynamic behavior. It is analogous to that of the Schwarzschild-AdS black hole, with no critical points(by solving the Eq. (17)) or phase transition phenomena present.
Reentrant phase transition regime ($0.3536<b<0.5$)  In this parameter range,by solving the Eq. (17), the system has two critical points and exhibits reentrant phase transition phenomena. As shown in the left diagram in Fig. 1, when the pressure is fixed at $P=0.98P_{z}$ ($P_{z}$ corresponds to the pressure at the onset of the zeroth-order phase transition), as the temperature decreases to $T_{1}$, the system undergoes a first-order phase transition from large black hole to small black hole. The Gibbs free energy of the two phases is continuous, but its first derivative is discontinuous, corresponding to the characteristics of a standard first-order phase transition. This phase transition process is depicted by a orange solid line in the right diagram in Fig. 1, showing a typical gas-liquid-type coexistence curve and ending at a critical point 1 $(P_{\rm c}^{(1)},T_{\rm c}^{(1)})$, where the heat capacity $C_{P}$ and isothermal compressibility $\kappa_{T}$ exhibit divergent behavior. When the temperature is further lowered to $T_{0}<T_{1}$, the system undergoes a transition from small black hole to large black hole, with the Gibbs free energy exhibiting a discontinuous jump $\Delta G\neq 0$ at temperature $T_{0}$, marking the occurrence of a zeroth-order phase transition. This zeroth-order phase transition starts at the point $(P_{z},T_{z})$ and ends at the triple point $(P_{t},T_{t})$. This phase transition region is extremely narrow, and we have determined its corresponding red solid line in the right diagram in Fig. 1. It is worth noting that this red line is not a traditional phase coexistence curve. Along this line, small black holes have a lower Gibbs free energy and remain in a locally stable state rather than coexisting phases; accordingly, the endpoint of this curve does not satisfy the critical point criterion. In a word, the first-order phase transition at high temperatures and the zeroth-order phase transition at low temperatures together form a complete reentrant phase diagram, profoundly reflecting the nontrivial coupling effects of BI nonlinear electrodynamics with the AdS background Rasheed:1997ns .
@@
Standard first-order phase transition regime ($b>0.5$)  In this parameter range, nonlinear effects become very weak, and the system behavior approaches that of linear Maxwell theory. As shown in Fig. 2,by solving the Eq. (17), the system has only one physical critical point. The the left diagram in Fig. 2 exhibits standard swallowtail behavior when the temperature decreases to $T_{2}$, the black hole undergoes a phase transition from small black hole to large black hole, displaying the typical first-order phase transition behavior analogous to that of four-dimensional RN-AdS black holes Kubiznak:2012wp .
@@ III Complex Phase diagram and Widom line
$2-2b^{2}r_{h}^{4}-\left(r_{h}^{2}-2b^{2}r_{h}^{4}-8\pi Pr_{h}^{4}\right)A=0.$ (18)